• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name peter

    Log entry time 08:02:48 on October24,2002

    Entry number 88938

    keyword=shift summary

    00:00 Shift begins with run # 2717 in progress.

    00:20 MCC called to say they are taking beam away for a few minutes to do a QE measurement.

    06:18 High voltage problem in the left arm - all the high voltages displayed seemed to have tripped, although the power supply was still 'ON'. We turned the HV off and then back on again - eventually everything went back to their normal operating parameters. Lost 8 minutes of data taking capability.

    06:30 Left arm Q1 crashed. The instructions to reset it remotely have disappeared from the counting house. After talking to the run coordinator, we have paged the tech on-call for assistance.

    07:56 Mark Stevens was unable to reset the Q1 remotely and so we have asked for a controlled access to reset the power supply.

    Run History - Q3D_d40 Kinematics

    Run # 2717 Good; 1M events; average current 95uA.
    Run # 2718 Good; 1M events; average current 101uA.
    Run # 2719 Good; 1M events; average current 93uA.
    Run # 2720 Good; 1M events; average current 92uA.
    Run # 2721 Good; 1M events; average current 84uA.
    Run # 2722 Good; 1M events; average current 98uA.
    Run # 2723 Junk; TS0 not responding, so run aborted.
    Run # 2724 Scintillator efficiency run.
    Run # 2725 Stopped early due to high voltage problem in left arm.
    Run # 2726 Stopped early due to left Q1 crash.